OpenStudy (anonymous):

In a triangle, the measure of the first angle is 24 degrees more than the measure of the second angle. The measure of the third angle is four times the measure of the second angle. If the total number of degrees in a triangle is 180, what is the measure of the largest angle?

OpenStudy (kkutie7):

Total degrees in an angle=180 a+b+c=180 a=24+b c=4*b this is what we know

OpenStudy (kkutie7):

180=(24+b)+b+(4*b) 180=24+b+b+4b 180=24+6b 156=6b 26=b

OpenStudy (kkutie7):

a=24+b a=24+26 a=50

OpenStudy (kkutie7):

c=4*b c=4*26 c=104 there ya go

OpenStudy (anonymous):

oh my gosh thanks!!!!!

OpenStudy (kkutie7):

any time =)